Sam was driving to Chicago. His rate for the first 80 miles was 10 mph faster than his rate for the last 140 miles. If his entire trip took 4 hours, what was his speed on each part of the trip? Round your answer to the nearest miles per hour
Answer by stanbon(75887)

Sam was driving to Chicago. His rate for the first 80 miles was 10 mph faster than his rate for the last 140 miles. If his entire trip took 4 hours, what was his speed on each part of the trip? Round your answer to the nearest miles per hour
----
First 80 mile DATA:
dist = 80 miles ; rate = r+10 mph ; time = 80/(r+10) hrs
--------------
Last 140 miles DATA:
dist = 140 miles ; rate = r mph ; time = 140/r hrs.
--------------------
Equation:
time + time = 4 hrs.
80/(r+10) + 140/r = 4 hrs
-----------------
80r + 140r + 1400 = 4r(r+10)
-----
20r + 35r + 350 = r^2+10r
r^2 - 45r - 350 = 0
-----------------------
r = 51.76 mph
-------------
Cheers,
Stan H.
-------